FEL Upgrade Project Weekly Brief
July 9-13, 2001

We achieved a new average power record for the FEL during the past two weeks: 2.1 kW- an increase of 24% over the previous record of 1.7 kW.  The FEL was operated with an wiggler efficiency  of 1%.  For the record the achieved power level and wiggler efficiency are now twice the original design specification.   In addition we tested FEL operation with the gun operated at the peak charges (135 pC) needed for the 10 kW Upgrade.  Details of both experiments are given in the Operations Section.

We received funding authorization ($404k) from ONR and DOE for the JTO contract for laser materials damage (i.e., lethality) studies. We thank all involved at JTO, ONR and the DOE JLab Site Office for blazing the path through the bureaucracies.

We received a summary report from SURA’s Maritime Technology Advisory Committee (MTAC) who reviewed our Navy funded FEL program on June 26.  MTAC graded our plan for a 100 kW Upgrade and felt the plan was credible and made several recommendations for strengthening our relationship with the Navy.

WBS 3 (Beam Physics):
Two days were spent last week doing orbit studies to benchmark our model of the lattice and optimize the beam transport for this week’s high power test.

WBS 4 (Injector):
Ran an additional curve on the N2 implanted faraday shielded Ti electrode at a 5 mm gap this time. The performance was worse than 1um polished Ti. The spacer nipple( 10-6 MV/m) for the gun has been manufactured.  We performed a dimensional stack up check on the gun drawings to confirm the proper anode to cathode gap was kept.

WBS 8 (Instrumentation):
The 3"  beam current cavity design is complete and an order is being placed for 5 pieces. The pulse truncator was modified and is awaiting further testing. This device allows for 0.5 microsecond increments in the pulse width for the FEL laser pulse. A prototype of a new BPM electronics is being worked on, testing should start in a couple weeks. Thanks to the EID group for there continued support.
Rich has been attending the Embedded Systems conference in Chicago this week.  Prototype boards were ordered for the new VME Timing Board. Cables were pulled and the new LINAC Patch Panels installed.  James built two new Pyro-electric Beam Position Monitors for testing. He and Dan also corrected a wiring problem with the Lab Status Box in User Lab 2 and wired Joe's Mirror Test Stand, also in Lab 2. Two user cameras were repaired, and 6 more ordered.  Dan built two Temperature Monitoring Chassis for RTD readout.
Patrick has the filter wheel working using the Rabbit TCP/IP interface and says the Optical Power Meter is coming along fine. This will allow for rapid connection of serial devices to the network and remotely monitor and control those devices without the FEL Users logging into EPICs.
Mike updated the search feature of the task system by offering a choice of the last 10 entries made for a particular task, or listing all entries pertaining to that task.  He also added the VME card inventory to Laser and condensed some data bases into a more useable format.
Drawings at EECAD: Analog/Differential Driver Board Schematic; Pyro-Electric Detector Buffer Amplifier Board Schematic, Fabrication, Assembly and Artwork; 32 Channel Sample and Hold Board Schematic, Assembly, Fabrication and Artwork; Ion Pump Power Supply Controller Schematic; User Lab Shutter Control Box Wiring, Schematic, Assembly and Fabrication; MPS System Drawing; Picomotor Relay Chassis Wiring Diagram; Generic Temperature Monitoring Chassis Schematic, Wiring, Assembly and Fabrication; OBPM/Picomotor Test Box
Wiring, Schematic, Assembly and Fabrication.

Drawings Completed:  LSS PLC Test Box, Fabrication, Wiring and Assembly; Beamviewer Chassis Test Box, Fabrication, Wiring and Assembly; Revisions to the LSS Master Control Chassis and User Lab Interface Box Wiring Diagrams. Check Plots Received: Beam viewer Interlock Box Wiring Diagram; LSS User Lab Interface Box Assembly Details; Molectron OBPM Schematic, Assembly, Fabrication and Artwork; Absolute Value/Sample and Hold Board Schematic, Assembly, Fabrication and Artwork.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  We will initiate a separate procurement for these magnets now and not wait for the completion of the design
    of the 180-degree dipoles (GY) as originally planned.  This package is complete while the GY requires some
    final adjustments to the model followed by the final tweaks to many of the details.
o  The coil, core and parts and the assembly specifications for the GW were written and the drawings printed for
    the sign off process.  Those specifications will serve as a template for the future dipole procurements.
o  We wrote the requisition and are starting discussions with Procurement to insure a smooth procurement
    process for the several batches of dipoles we will need.
Injector Dipoles (DU/DV)
o  The DULY continues working on the revisions to the drawings of the Injector Dipoles (GU & GV).
Arc Dipoles (GY, GX, GQ)
o  AES continued design of the Reverse Bend (GQ).  Core pieces were detailed and coil assembly ad details are
    being worked on.
QX (3 inch quad)
o  I mistakenly reported that the core contract was placed in the last report.  Rather, some additional information
    was required of the vendors.  The recommendation for award is now under internal review.
o  We made some progress and gained insight in eliminating higher order multipoles due the pole tips’ end field
    roll off.  We brought the sum of the multipoles on the Prototype from 0.62 % to 0.42 with more progress
    expected.  This will be a last tweak to the core drawings that we can introduce after the core order is placed.
QT (Trim Quad)
o  The magnetic model is acceptable.  Most of the design and details are drawn.  When the designer returns from
    vacation next week, the change of the details to match the final model will start.
o  Robin Wines has not been able to continue with magnetic modeling over the last reporting period because of
    the octupole work below and her duties as magnet system owner for CEBAF had her looking at cooling
    problems with the septa magnets for much of her time.
o  The task order for DULY Research to start magnetic modeling and design was sent to Procurement.
o  The Dipole Steel order was placed with Bethlehem Lukens Steel with a 10-week delivery.
Beam Line and Vacuum
o  Work continued on the Interface Control Document for the task definition for the Arc and Optical Chicane
o  The review period for the set of the drawing of the girders for the return leg was completed.
o  Work continued on the drawing set for the girders between Cryomodules.
o  Work started on the layout of the injection and recombination area.

WBS 11 (Optics):
Detailing continues on the internal mounting hardware in the optical cavity chambers.  We received drawings of the mirror design employed by the AVLIS, and now NIF (courtesy of R. Chow - LLNL), and are evaluating their applicability.

We decided to abandon (at least for now) use of two insertable viewers to route beam to the diagnostics mounted between the outcoupler chamber and the collimator.  The designer found that dual-action air cylinders exist and should do the job.

Our laser interferometer was installed by the vendor's application engineer.  This will be used in general to confirm optics' specifications, and specifically, with the mirror test stand.

The new drive laser specifications and objectives was turned over to Procurement.

Support for operations/experimenters:
 Installed steering and condensing optics to mirror test stand.  Installed fast IR detector to checkout Drive Laser Intermediate Controller DLIC (built by the I&C group).  Provided and installed remote-controlled 5-axis stage for X-ray experiment.  Repaired output mirror on FIR spectrometer (for JLab THz experiment). Provided optics hardware and support to A. Reilly and M. Kelley (both at College of William and Mary).  Provided support to JTO experiment in the form of beam profile calibration and sample documentation (using I & C's incredible digital
camera!).  Installed power meter for FEL efficiency measurements.  Installed new lamps in drive laser.  Changed mirrors in order to shift operation from 3 microns to 6 microns.  Inspected recoated, broadband cavity mirrors on ZnSe substrates.  They passed this time.  Received new optics to support third harmonic lasing at 1.6 microns.  They will be checked during the next week.

Ops ran smoothly over the week. The first part of the week was devoted measuring ablation at 2.9 microns.   That was followed by X-ray Thompson scattering measurements.  The X-ray measurements successfully mapped the distribution of scattering of 3 micron light for the first time in preparation for a double scattering experiment to be performed in the fall.  That effort was followed by a high power demonstration run. We were repeatedly able to run to 2.1 kW using 65 pC for 0.97% extraction efficiency, a 24% improvement over our previous CW efficiency.

We then re-cesiated the cathode, this being the first re-cesiation since Feb.16.   The QE improved from 0.5% to 1.9% and we again verified the high power FEL performance after the recesiation.  Since average current was still limited to 5 mA no significant additional power was produced.  We did not take the additional time it would take to optimize the setup for high charge running at 135 pC although we were able to lase strongly there with 1% pulsed efficiency.  That effort was followed by a switch over to 6 micron mirrors and we brought up the system and established lasing for user tests at 5.8 microns for next week.

We also were able to see and measure the spectrum of the collective far IR radiation emitted from the electron bunches in going around our bends.  It is expected that this radiation will be useful as a beam diagnostic and a research tool in its own right.

Three runs in an entirely parasitic mode were accomplished for the nanotube collaboration (CWM, NASA, PSU).  They are continuing to collect nanotubes at various target/laser configurations to optimize the production rate of single wall tubes.